OpenStudy (anonymous):

limit as x approaches 0 for (cubed root(125 + x)) -5 / (x)

OpenStudy (anonymous):

Apply L'Hospital Rule.

OpenStudy (anonymous):

\[lim_{x\to0}\left(\frac{\sqrt[3]{125+x}-5}{x}\right)\] We can apply the L'Hopital Rule. \[lim_{x\to0}\frac{f(x)}{g(x)}=lim_{x\to0}\frac{f'(x)}{g'(x)}\] then \[lim_{x\to0}\left(\frac{1}{3*\sqrt[3]{(125+x)^2}}\right)\] replace the approache \[lim_{x\to0}\left(\frac{1}{3*\sqrt[3]{(125+0)^2}}\right)=\frac{1}{3*\sqrt[3]{125}}=\frac{1}{15}\]